# EvidenceMap: Learning Evidence Analysis to Unleash the Power of Small Language Models for Biomedical Question Answering

### v2

@@ Abstract
When addressing professional questions in the biomedical domain, humans typically acquire multiple pieces of information as evidence and engage in multifaceted evidence analysis to provide high-quality answers. Current LLM-based answer generation methods lack a detailed definition and learning process for evidence analysis, leading to the risk of error propagation and hallucinations while using evidence. Although increasing the parameter size of LLMs can alleviate these issues, it also presents challenges in model training and deployment with limited resources. In this study, we propose EvidenceMap, which aims to enable a tiny pre-trained language model to explicitly learn multiple aspects of biomedical evidence, including supportive evaluation, logical correlation and content summarization, thereby latently guiding a small generative model (around 3B parameters) to provide textual responses. Experimental results demonstrate that our method, fine-tuning a language model with 66M parameters, exceeds the RAG method with an 8B LLM by 19.9% and 5.7% in reference-based quality and accuracy, respectively.
@@
Chang Zong1, Jian Wan1, Siliang Tang2, Lei Zhang1 thanks: Corresponding author. 1Zhejiang University of Science and Technology 2Zhejiang University zongchang@zust.edu.cn
@@ 1 Introduction
Answering specialized questions typically requires gathering a wide range of information as evidence, such as relevant texts from academic articles and domain knowledge Louis et al. (2024); Singhal et al. (2025). In the biomedical domain, it is crucial to formulate answers by evaluating and summarizing various evidence Singhal et al. (2023); Zhang et al. (2024). Figure 1 depicts a modern scenario in which a biomedical expert acquires relevant evidence from PubMed papers and a large language model (LLM) and performs a series of analytical steps that include supportive evaluation Jin et al. (2021), logical correlation Chen et al. (2021), and content summarization Xie et al. (2022), ultimately leading to a high-quality answer.
Figure 1: A diagram illustrates the evidence analysis process when addressing biomedical questions.
Generative question answering based on LLM has shown advantages in recent years. Current methods primarily leverage external knowledge for implicit reasoning Lewis et al. (2020); Asai et al. (2023); He et al. (2024); Wang et al. (2024) or utilize the ability of LLMs to generate explicit thinking processes Wei et al. (2022); Lightman et al. (2023); Yao et al. (2024), thus mitigating hallucinations and producing accurate and comprehensive responses. However, as illustrated in Figure 2, these approaches still encounter error propagation when confronted with various evidence due to lack of explicit analysis learning. Moreover, although increasing the scale of model parameters can improve reasoning capabilities, it poses challenges for training and deployment under limited resources.
Figure 2: Existing methods (left) suffer from unreasonable use or incorrect reasoning on low-quality evidence. Our framework (right) reduces error propagation by learning evidence analysis with small language models.
To alleviate the aforementioned issues, this study focuses on answering biomedical questions with various evidence and proposes a novel framework, named EvidenceMap. The framework explicitly learns and utilizes evidence analysis through small language models, thereby efficiently mimicking human problem-solving processes. Specifically, our framework models an evidence map using multiple pieces of evidence relevant to the question, and subsequently employs a small pre-trained language model (PLM) to learn the representation of components in evidence analysis including supportive evaluation, logical correlation, and content summarization, which serve as the analytical prompt to drive a small generative language model (GLM) to autoregressively derive answers. The experimental results demonstrate that applying evidence analysis to small language models significantly improves biomedical question-answering performance.
Our contributions are summarized as follows.
@@
  • Extensive experiments and analyses on biomedical datasets demonstrate that our approach, by explicitly learning and utilizing evidence analysis, achieves better performance on smaller models compared to larger-scale reasoning methods.
@@ 2.1 Evidence Map Modeling
We construct an evidence map ($\mathcal{M}$) for each input question to facilitate reasoning during generation. An evidence map introduces three key types of information, including the evaluation of how the evidence supports the question ($\mathcal{R}_{eval}$), the correlation between the evidence ($\mathcal{R}_{cor}$), and the summarization of all the evidence ($E_{sum}$). Generally, an evidence map can be formulated as follows.
@@ 2.2 Learning Generative QA with LMs
$A^{*}=f_{glm}\big{(}Q,f_{plm}(\theta^{*}),\mathcal{M}\big{)},$ (4)
where $f_{plm}(\theta^{*})$ is a tiny pre-trained model (e.g. DistilBERT-Base), $f_{glm}$ is a small generative model (e.g. Llama-3.2-3B) for providing the best answer $A^{*}$. Then, our learning process is to find the optimized parameter set $\theta^{*}$, which leads to a minimum generative loss for all question-answer pairs, given the corresponding evidence map $EM_{i}$ for each question $Q_{i}$. This can be denoted as below.
$\theta^{*}=\mathop{\arg\min}\limits_{\theta}\sum_{i}\mathcal{L}\Big{(}A_{i},f_{glm}\big{(}Q_{i},f_{plm}(\theta),\mathcal{M}_{i}\big{)}\Big{)}.$ (5)
Figure 3: The overview of our EvidenceMap framework. An evidence-augmented question is modeled as an evidence map, in which the representations of evidence evaluation, correlation and summarization are computed and updated by a tiny pre-trained model accordingly, in order to facilitate answering by a small generative model.
@@ 3.1 Evidence Acquisition
As described in Figure 1, we gather relevant evidence from multiple sources for each question, including the internal knowledge of an LLM and the external information from academic papers. Specifically, for the evidence of an LLM, we ask GPT-4o Hurst et al. (2024) to take advantage of its knowledge to provide a series of key points that support the question, described in a concise language (see Appendix A.2). In addition, we extract paper snippets from existing biomedical QA datasets such as BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019), which have already been annotated by human experts, as evidence related to a question. Therefore, an evidence-augmented QA dataset is constructed, in which each question corresponds to one piece of LLM evidence $E_{lm}$ and multiple pieces of paper evidence from $E_{p}^{1}$ to $E_{p}^{n}$. The input data can be described as follows.
@@ 3.3 Analysis-Augmented Generation
With the result of the analysis by handling an evidence map, the final answer to the question is provided in an autoregressive way using a generative language model $f_{glm}$. We first convert all textual evidence and the question into an embedding $H_{t}$ using a text embedder, which is the first embedding layer of $f_{glm}$ as follows.
@@
Then the evidence analysis results are projected into the same space with the textual embedding using a linear layer, as a soft prompt as follows.
@@
where $\mathcal{R}_{eval}$ and $\mathcal{R}_{cor}$ are the sets of all the supportive relationships $R_{eval}^{i}$ and logical correlations $R_{cor}^{(i,j)}$, respectively. We then prepend the analysis embedding $\hat{H}_{a}$ to the textual embedding $H_{t}$ as the overall prompt to implement an analysis-augmented generation with the frozen language model $f_{glm}$, which is formulated as below.
@@ 3.4 Training Objective
The autoregressive training objective concentrates on updating the parameter set $\theta$ of a small language model to provide the representation of evidence analysis to predict subsequent tokens accurately by a generative model. We calculate the probability of generating the $i$-th token of the target answer $A$ with the overall prompt and all previous tokens as follows.
$\mathcal{L}=\sum_{i=1}^{L}log\big{(}A_{i}|\texttt{Prompt}_{all},A_{0:i-1};\theta\big{)}.$ (14)
@@ 4.3 Main Results
In order to demonstrate the effectiveness of learning evidence analysis, we consider three configurations: 1) Inference-LLM: Using larger language models in general and biomedical domains, including BioMed-Llama Pal and Sankarasubbu (2024), Llama-3.1 Dubey et al. (2024), and GPT-4o-mini Hurst et al. (2024). 2) Inference-SLM: Using a smaller frozen generative model with different reasoning strategies, such as RAG Lewis et al. (2020); Gao et al. (2023), CoT Wei et al. (2022) and RAT Wang et al. (2024); 3) Tuning-SLM: Adapting the task by fine-tuning a generative model with LoRA Hu et al. (2021) or updating the pre-trained SLM via a prompting tuning style Lester et al. (2021); Li and Liang (2021). More details about baseline settings are provided in Appendix B.
@@
The results of the performance comparison are shown in Table 1. We can make the following observations:
First, Learning evidence analysis significantly improves the ability to utilize diverse information. When using the same generative language model (Llama-3.2-3B), EvidenceMap outperforms RAG, CoT and RAT, demonstrating that our approach effectively leverages the relationships among various pieces of evidence and between evidence and the question, thereby maximizing the value of multifaceted information.
Second, Learning evidence analysis unleashes the power of small language models to efficiently resolve biomedical questions. Compared to larger models, including the RAG method with general LLMs (GPT-4o-mini, Llama-3.1-8B) and biomedical LLM (OpenBioLLM-8B), our approach demonstrates superior performance even with a 1B model. This indicates that evidence analysis can effectively unleash the capabilities of small language models.
@@ 4.4 Ablation Study on Evidence Analysis
Table 2: Ablation study of three parts within the evidence analysis on two datasets. w/o Eval, w/o Cor, and w/o Sum represent removing the evaluation, correlation, and summarization, respectively.
As shown in Table 2, each part contributes to the overall performance, which is consistent with our hypothesis. In detail, the logical correlation has the most significant impact on overall performance on both datasets, indicating that determining the logical relationship between two pieces of evidence is crucial for providing a high-quality answer. Meanwhile, evaluating how evidence supports the question plays a more significant role than summarizing the evidence, indicating that assessing the relationship between evidence and the question improves effectiveness. Furthermore, incorporating evidence analysis has a greater effect on the BioASQ dataset compared to the PubMedQA dataset in general, showing that the effect of evidence analysis is influenced by specific domains.
@@ 4.5 Ablation Study on Answer Generation
We further analyze the impact of each module in the answer generation step. We denote the method that does not incorporate textual evidence embeddings as "w/o TE", and the one that eliminates the input of evidence analysis as "w/o EA". We replace the MLP projector with a simple linear layer and denote it as "w/o Proj". Performance results and variations are illustrated in Table 3. The results indicate that both the textual input and the analysis of the evidence have a significant impact on the final performance, which can significantly improve the quality of answers beyond the foundation of basic language model generation. In contrast, the projector used to align the evidence analysis results with the generative language model is just an auxiliary to the overall performance.
@@
Table 3: Ablation study on the step of answer generation on two biomedical datasets. TE, EA, and Proj are modules from the step representing textual embedding, evidence analysis, and projector, respectively.
@@ 4.6 Impact of Pre-trained Language Models
More powerful pre-trained language models can achieve better results. BERT-Base, GPT2, and ModernBERT, which have larger parameter sizes and better performance in general language abilities, improve results on both datasets. This indicates that as the language model continues to update and advance, the capabilities of our framework will also be further enhanced.
@@ 5 Case Study
In this section, we qualitatively analyze the effectiveness of our framework in addressing biomedical questions by presenting specific cases. Figure 5 presents two examples of utilizing evidence in distinct ways to formulate responses. In each case, a question is provided alongside relevant evidence, the standard answer, and the answers generated by RAG and our EvidenceMap.
Case 1 illustrates a scenario of factual questions. We notice that the content of the standard answer is consistent with the evidence presented in the paper, both referencing the drug combination "Pembrolizumab plus lenvatinib", which is more accurate and comprehensive than the LLM evidence. The answer provided by EvidenceMap is incomplete, mentioning only the drug "Pembrolizumab". In contrast, the response generated by the RAG method references "T-VEC", which is an entirely different medication, which results in the dissemination of incorrect information and leads to a decrease in accuracy. This indicates that although our framework may provide incomplete information, analyzing the relationships between pieces of evidence can mitigate the hallucination problem in generative models and help prevent the generation of incorrect answers.
Case 2 illustrates a scenario that addresses an explanatory question. By analyzing the standard answers, we can tell that a more complex analytical process is necessary. Accurately addressing this question requires reliance on information from a piece of paper evidence and the LLM evidence, while simultaneously avoiding certain elements from other papers and the misleading part of the LLM evidence. Specifically, the "IL-12 family" mentioned in the standard answer should be derived from the LLM evidence, which points out the "interleukin-12 cytokine family". Our EvidenceMap accurately captures this information element and presents it in the response. Meanwhile, the standard answer mentions "immunoreaction and promoting endometrial cell proliferation" a concept supported by evidence from the second paper. EvidenceMap precisely identifies this content and incorporates it into the response. In contrast, the answer generated by the RAG method overlooks these key pieces of information and utilizes extraneous content ("T cells") from other papers. This case illustrates that our approach effectively utilizes multiple and diverse sources of evidence by analyzing the relationships between them, thereby providing more accurate answers.
@@ 6.1 Biomedical Evidence Analysis
Resolving issues in the biomedical domain requires a multifaceted analysis of related evidence. A recent study Singhal et al. (2025) suggests that addressing biomedical questions requires correct retrieval, accurate manipulation, and complete understanding of relevant knowledge. Some studies Jin et al. (2021); Singhal et al. (2023) indicate that models need to ignore irrelevant distractors to avoid incorrect interpretations of the evidence in the answers. Other studies focus on logical reasoning Chen et al. (2021) and summarization of evidence Xie et al. (2022); Zhang et al. (2024), thereby integrating medical evidence to support decision analysis and provide high-quality answers. Our study builds on these foundational works, integrating the above key aspects of evidence analysis, and propose a framework for evidence analysis and analysis-augmented generation to address biomedical question-answering tasks.

### v4

@@ Abstract
When addressing professional questions in the biomedical domain, humans typically acquire multiple pieces of information as evidence and engage in multifaceted analysis to provide high-quality answers. Current LLM-based question answering methods lack a detailed definition and learning process for evidence analysis, leading to the risk of error propagation and hallucinations while using evidence. Although increasing the parameter size of LLMs can alleviate these issues, it also presents challenges in training and deployment with limited resources. In this study, we propose EvidenceMap, which aims to enable a tiny pre-trained language model to explicitly learn multiple aspects of biomedical evidence, including supportive evaluation, logical correlation and content summarization, thereby latently guiding a small generative model (around 3B parameters) to provide textual responses. Experimental results demonstrate that our method, learning evidence analysis by fine-tuning a model with only 66M parameters, exceeds the RAG method with an 8B LLM by 19.9% and 5.7% in reference-based quality and accuracy, respectively.
@@
Chang Zong1, Jian Wan1, Siliang Tang2, Lei Zhang1 1Zhejiang University of Science and Technology 2Zhejiang University zongchang@zust.edu.cn
@@ 1 Introduction
Answering professional questions typically requires gathering a wide range of information as evidence, such as relevant academic articles and domain knowledge from experts Louis et al. (2024); Singhal et al. (2025). In the biomedical domain, it is crucial to formulate answers by evaluating and summarizing various evidence Singhal et al. (2023); Zhang et al. (2024). We describe a modern scenario in Figure 1, in which a biomedical expert acquires relevant evidence from PubMed papers and a large language model (LLM) and performs a series of analytical steps that include supportive evaluation Jin et al. (2021), logical correlation Chen et al. (2021), and content summarization Xie et al. (2022), ultimately delivering a high quality answer.
Figure 1: A diagram describing the analysis with various evidence when addressing biomedical questions.
To answer domain questions, LLM-based generation methods have shown advantages in recent years. Current methods primarily incorporate external knowledge into LLMs for reasoning Lewis et al. (2020); Asai et al. (2023); He et al. (2024); Wang et al. (2024) or enable LLMs to generate thinking processes Wei et al. (2022); Lightman et al. (2023); Yao et al. (2024), thus mitigating hallucinations and producing accurate and comprehensive responses. However, as illustrated in Figure 2, due to a lack of training in analysis skills and descriptions of evidence interactions, these approaches still face the issue of error propagation when addressing biomedical questions. Moreover, although increasing the scale of model parameters can improve reasoning capabilities, it poses challenges for training and deployment under limited resources.
Figure 2: Existing methods (left) suffer from erroneous thinking process or reasoning on low-quality evidence. Our framework (right) reduces error propagation by learning evidence analysis explicitly and efficiently with small language models (SLM).
To alleviate the aforementioned issues, this study focuses on biomedical question answering with various evidence, and proposes a novel framework, named EvidenceMap. The framework explicitly learns evidence analysis with small language models, thereby efficiently mimicking the human-solving process in the biomedical domain. Specifically, our framework defines a conceptual model called Evidence Map, which describes the relational dynamics among multiple pieces of evidence relevant to the question, and subsequently employs a small pre-trained language model (PLM) to learn the representation of analytical elements in the Evidence Map, including supportive evaluation, logical correlation, and content summarization, which serve as the prompt to drive a small generative language model (GLM) to produce answers. The experimental results indicate that learning evidence analysis can significantly enhance the performance of biomedical question answering. Our contributions are summarized as follows.
@@
  • Extensive experiments and analyses on biomedical datasets demonstrate that our approach, by explicitly learning and utilizing evidence analysis, can significantly improve the quality of answers and outperforms reasoning methods with larger-scale language models.
@@ 2.1 Evidence Map Modeling
We define an evidence map ($\mathcal{M}$) for each input question to facilitate reasoning during generation. An evidence map introduces three key types of information, including the evaluation of how the evidence supports the question ($\mathcal{R}_{eval}$), the correlation between the evidence ($\mathcal{R}_{cor}$), and the summarization of all the evidence ($E_{sum}$). Generally, an evidence map can be formulated as follows.
@@ 2.2 Learning Generative QA with LMs
$A^{*}=f_{glm}\big{(}Q,f_{plm}(\mathcal{M};\theta^{*})\big{)},$ (4)
where $f_{plm}(\cdot)$ is a tiny pre-trained model (e.g. DistilBERT-Base), $f_{glm}(\cdot)$ is a small generative model (e.g. Llama-3.2-3B) for providing the best answer $A^{*}$. Then, our learning process is to find the optimized parameter set $\theta^{*}$, which leads to a minimum generative loss for all question-answer pairs, given the corresponding evidence map $\mathcal{M}_{i}$ for each question $Q_{i}$. This can be denoted as below.
$\theta^{*}=\mathop{\arg\min}\limits_{\theta}\sum_{i}\mathcal{L}\Big{(}A_{i},f_{glm}\big{(}Q_{i},f_{plm}(\mathcal{M}_{i};\theta)\big{)}\Big{)}.$ (5)
Figure 3: The overview of our EvidenceMap framework. An evidence-augmented question is modeled as an evidence map, in which the representations of evidence evaluation, correlation and summarization are computed by a pre-trained language model (PLM) accordingly, in order to facilitate answering by a generative model (GLM).
@@ 3.1 Evidence Acquisition
As described in Figure 1, we gather relevant evidence from multiple sources, including a powerful commercial LLM and academic papers, for each question. Specifically, for evidence from an LLM, we ask GPT-4o Hurst et al. (2024), a powerful commercial LLM, to utilize its own knowledge to provide a series of key points that support the question, which are described concisely (see Appendix A.2). In addition, we directly extract paper snippets from existing biomedical QA datasets such as BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019), which have already been annotated by human experts, as the evidence related to the question. Therefore, an evidence-augmented QA dataset is constructed, in which each question corresponds to one piece of LLM evidence $E_{lm}$ and multiple pieces of paper evidence from $E_{p}^{1}$ to $E_{p}^{n}$. The input data can be described as follows.
@@ 3.3 Analysis-Augmented Generation
With the result of the analysis from an evidence map, the final answer to the question is provided in an autoregressive way using a generative language model $f_{glm}$. We first convert all textual evidence and the question into an embedding $H_{t}$ using the first embedding layer of $f_{glm}$ as follows.
@@
Then the analysis results are projected into the same space as the textual embedding to serve as a soft prompt, using a MLP network with each linear layer formualted as below.
@@
where $R_{eval}^{i}\in\mathcal{R}_{eval}$ and $R_{cor}^{(i,j)}\in\mathcal{R}_{cor}$ are the sets of all the supportive and logical relationships, respectively. We then prepend the analysis embedding $\hat{H}_{a}$ to the textual embedding $H_{t}$ as the overall prompt to implement an analysis-augmented generation with the frozen language model $f_{glm}$, which is formulated as below.
@@ 3.4 Training Objective
The autoregressive training objective concentrates on updating the set of parameters $\theta$ of a small pre-trained language model to provide the representation of evidence analysis to predict subsequent tokens accurately using a generative model. We calculate the loss with the probability of generating the $i$-th token of the target answer $A$ with the overall prompt and all previous tokens as follows.
$\mathcal{L}=\sum_{i=1}^{L}log\big{(}A_{i}|\texttt{Prompt}_{all},A_{0:i-1};\theta\big{)},$ (14)
where $L$ is the length of the target answer $A$.
@@ 4.3 Main Results
In order to demonstrate the effectiveness of learning evidence analysis, we consider three configurations: 1) Inference-LLM: Using larger language models in general and biomedical domains, including Llama-3.1 Dubey et al. (2024), GPT-4o-mini Hurst et al. (2024), and OpenBioLLM Pal and Sankarasubbu (2024). 2) Inference-SLM: Using a smaller frozen generative model with various reasoning strategies, such as RAG Lewis et al. (2020); Gao et al. (2023), CoT Wei et al. (2022) and RAT Wang et al. (2024); 3) Tuning-SLM: Adapting the task by fine-tuning a small generative model with LoRA Hu et al. (2021) or updating a small pre-trained model via a soft prompt tuning style Lester et al. (2021); Li and Liang (2021). More details on baseline settings are provided in Appendix B.
@@
The results of the performance comparison are shown in Table 1. We can make the following observations.
First, Learning evidence analysis significantly improves the ability to utilize diverse information. When using the same generative language model (Llama-3.2-3B), EvidenceMap outperforms RAG, CoT and RAT, demonstrating that our approach effectively leverages the relationships among multiple pieces of evidence and the question, thereby demonstrating the advantages of multifaceted analysis.
Second, Learning evidence analysis unleashes the power of small language models to resolve biomedical questions. Compared to larger models, including the RAG method with general LLMs (GPT-4o-mini, Llama-3.1-8B) and a biomedical LLM (OpenBioLLM-8B), our approach demonstrates superior performance even with a 1B model, indicating that evidence analysis can effectively unleash the ability of smaller language models.
@@ 4.4 Ablation Study on Evidence Analysis
Table 2: Ablation study of three aspects within the evidence analysis on two datasets. The results indicate that all aspects of the analysis we proposed have positive impacts on the performance.
As shown in Table 2, each part contributes to the overall performance, which is consistent with our hypothesis. In detail, the logical correlation has the most significant impact on overall performance on both datasets, indicating that determining the logical relationship between two pieces of evidence is crucial for providing a high-quality answer. Meanwhile, evaluating how evidence supports the question plays a more significant role than summarizing the evidence, indicating that assessing the relationship between evidence and the question improves effectiveness. Furthermore, incorporating evidence analysis has a greater effect on the BioASQ dataset compared to the PubMedQA dataset in general, showing that the effect of evidence analysis is influenced by the specific data distribution.
@@ 4.5 Ablation Study on Answer Generation
We further analyze the impact of each module in the answer generation step. We denote the method that does not incorporate textual evidence embeddings as "w/o TE", and the one that eliminates the input of evidence analysis as "w/o EA". We replace the MLP projector with a simple linear layer and denote it as "w/o Proj". Performance results and variations are illustrated in Table 3. The results indicate that both the textual input and our proposed evidence analysis have significant impacts on the final performance, which can dramatically improve the quality of answers beyond the foundation of basic language model generation. In contrast, the projector used to align the evidence analysis results with the generative language model is just an auxiliary to the overall performance.
@@
Table 3: Ablation study on the answer generation step. Our proposed evidence analysis (EA) demonstrates a significant impact on the effectiveness of answer generation, which is secondary only to that of text input (TE).
@@ 4.6 Impact of Pre-trained Language Models
More powerful pre-trained language models can achieve better results. BERT-Base, GPT2, and ModernBERT, which have larger parameter sizes and better performance in general language abilities, improve results on most indicators. This indicates that as the language model continues to update and advance, the capabilities of our framework will also be further enhanced.
@@ 5 Case Study
In this section, we qualitatively analyze the effectiveness of our framework in addressing biomedical questions by presenting specific cases. Figure 5 presents two examples of utilizing evidence in distinct ways to formulate responses. In each case, a question is provided alongside the relevant evidence, the reference answer, and the answers generated by the 3B version of RAG and our EvidenceMap.
Case 1 illustrates a scenario of factual questions. We notice that the content of the reference answer is consistent with the evidence presented in the paper, both referencing the drug combination "Pembrolizumab plus lenvatinib", which is more accurate and comprehensive than the LLM evidence. The answer provided by EvidenceMap is incomplete, mentioning only the drug "Pembrolizumab". In contrast, the response generated by the RAG method references "T-VEC", which is an entirely different medication, which results in the dissemination of incorrect information and leads to a decrease in accuracy. This indicates that although our framework may provide incomplete information, analyzing the relationships between pieces of evidence can mitigate the hallucination problem in generative models and help prevent the generation of incorrect answers.
Case 2 illustrates a scenario that addresses an explanatory question. By analyzing the reference answers, we can tell that a more complex analytical process is necessary. Accurately addressing this question requires reliance on information from a piece of paper evidence and the LLM evidence, while simultaneously avoiding certain elements from other papers and the misleading part of the LLM evidence. Specifically, the "IL-12 family" mentioned in the reference answer should be derived from the LLM evidence, which points out the "interleukin-12 cytokine family". Our EvidenceMap accurately captures this information element and presents it in the response. Meanwhile, the reference answer mentions "immunoreaction and promoting endometrial cell proliferation" a concept supported by evidence from the second paper. EvidenceMap precisely identifies this content and incorporates it into the response. In contrast, the answer generated by the RAG method overlooks these key pieces of information and utilizes extraneous content ("T cells") from other papers. This case illustrates that our approach effectively utilizes multiple and diverse sources of evidence by analyzing the relationships between them, thereby providing more accurate answers.
@@ 6.1 Biomedical Evidence Analysis
Resolving biomedical questions requires a multifaceted analysis of related evidence. A recent study Singhal et al. (2025) suggests that addressing biomedical questions requires correct retrieval, accurate manipulation, and complete understanding of relevant knowledge. Some studies Jin et al. (2021); Singhal et al. (2023) indicate that models need to ignore irrelevant distractors to avoid incorrect interpretations of the evidence in the answers. Other studies focus on logical reasoning Chen et al. (2021); Kang et al. (2023) and summarization of evidence Xie et al. (2022); Zhang et al. (2024), thereby integrating medical evidence to support decision analysis and provide high-quality answers. Our study builds on these foundational works, integrating the above key aspects of evidence analysis, and proposes a framework for evidence analysis and analysis-augmented generation to address biomedical question-answering tasks.